# FUIA: Model Inversion Attack against Federated Unlearning

### v2

@@ Abstract
With the enactment of regulations related to the “right to be forgotten”, federated learning is facing new privacy compliance requirements. To meet these demands, researchers have proposed federated unlearning (FU) methods. However, existing FU studies primarily focus on improving unlearning efficiency, with little attention given to the potential privacy risks introduced by FU itself. To bridge this research gap, we propose a novel Federated Unlearning Inversion Attack (FUIA) to expose potential privacy leakage in FU methods. This work represents the first systematic study on the privacy vulnerabilities inherent in FU. FUIA applies to three major FU types: sample unlearning, client unlearning, and class unlearning, demonstrating broad applicability and threat potential. Specifically, in FUIA, the server acts as an honest-but-curious adversary, continuously recording model parameters throughout the unlearning process. Inspired by gradient inversion attacks in FL, we estimate the gradient information of forgotten data by analyzing the model parameter differences before and after unlearning. This enables us to reconstruct the features or labels of the forgotten data. The existence of FUIA directly contradicts the goal of FU, which is to eliminate the influence of specific data. Instead, it exploits vulnerabilities in the FU process to recover forgotten data, revealing significant privacy risks in current FU methods. To mitigate these risks, we explore two potential defense strategies that introduce a trade-off between privacy protection and model performance. Extensive experiments on multiple benchmark datasets and various FU methods demonstrate that FUIA effectively reveals private information of forgotten data.
@@
Gradient inversion attack, machine unlearning, federated learning.I Introduction
With the rapid development of artificial intelligence, the issue of data privacy and security in training machine learning models has garnered increasing attention from both academia and industry [1]. To address these concerns, various regulations have been enacted worldwide, such as the General Data Protection Regulation (GDPR) [2] in the European Union and the California Consumer Privacy Act (CCPA) [3] in the United States. These regulations provide robust legal protections for data owners, with the “right to be forgotten” [4] drawing significant attention in recent years. This right mandates that when a data owner requests the deletion of their data, the model owner must not only remove the corresponding data from the database but also eliminate its impact on the trained model. Federated Learning (FL) [5] is designed to enable distributed model training while preserving the privacy of individual clients. However, the introduction of the “right to be forgotten” imposes new requirements on FL, particularly in efficiently fulfilling data owners’ requests for data removal. Recent studies have proposed a novel approach known as Federated Unlearning (FU) [6], which aims to erase the influnence of the target data from the model without the need for retraining the entire model.
@@
During the FU process, the server has full access to the differences between the original model and the unlearned model, even down to the parameter updates uploaded by each client in every round of the unlearning process. If the server acts maliciously, this extensive knowledge could lead to severe privacy risks, particularly regarding the exposure of forgotten data. In FL, gradient inversion attack (GIA) is designed based on this exact principle, leveraging the observability of model updates to reconstruct the original training data of clients [15, 16]. Inspired by GIA, we propose FUIA, which aims to infer the features or other information of the forgotten data. The attack fundamentally contradicts the core objective of FU, which is to erase all traces of target forgotten data. In the FUIA, the attacker is defined as an “honest-but-curious” server, meaning it follows the prescribed training and unlearning protocols but continuously records all parameter updates from clients throughout the entire process. This strategy of storing parameter updates is widely employed in FL and FU privacy research and does not impose additional assumptions. As shown in Figure 1, the attacker in FUIA exploits the differences between the stored original model parameters and the unlearning model parameters to recover the features of the forgotten data or infer its deep semantic information.
@@
Based on different unlearning targets, existing research on FU can be classified into three categories: sample unlearning, client unlearning, and class unlearning [17]. Sample unlearning focuses on eliminating the influence of several data samples. Client unlearning requires the model to entirely forget the data from a specific client, making sure that the contributions of the client are no longer reflected in the model parameters. Class unlearning aims to completely remove the influence of a specific class of data from the model. To exploit vulnerabilities in these FU methods, we propose FUIA, which targets all three types of unlearning mechanisms. In the cases of sample unlearning and client unlearning, FUIA aims to recover the forgotten data information, such as reconstructing images that were intended to be removed. The core principle behind this attack lies in the observation that the parameter differences before and after unlearning inherently encode gradient information related to the forgotten data. By leveraging techniques inspired by GIA, which are well-studied in FL, the adversary can apply optimization-based inversion methods to recover the feature representation of the forgotten data. For class unlearning, since the server already possesses knowledge of the global dataset, the goal of FUIA shifts from reconstructing individual data samples to inferring the label of the forgotten class. In general, data from a particular class has a significant impact on model parameters, especially in the output layer. Therefore, the adversary can compare the parameter changes of the output layer before and after unlearning to infer the label of the removed class. The attack strategy of FUIA demonstrates its broad applicability across different FU scenarios. This finding highlights potential privacy risks in FU and poses a significant challenge to the current FU design, emphasizing the need for future research to develop more robust and secure unlearning mechanisms.
@@
To address the privacy challenges posed by FUIA, we draw inspiration from existing methods in FL designed to defend against GIA [18, 15]. Based on these insights, we propose two general defense strategies. First, during the unlearning process in FU, we introduce gradient pruning, where a certain proportion of less important parameter updates from each client is removed before aggregation. The primary goal of this method is to prevent the server from obtaining a complete and precise set of model updates. Second, we leverage differential privacy as an additional defense mechanism. During the unlearning process, we introduce random noise into the model updates submitted by clients, ensuring that the server only receives perturbed parameter updates. These approaches effectively both disrupt the ability of the adversaries to infer precise gradient information related to the forgotten data, thereby weakening the effectiveness of FUIA. However, they can lead to poor model performance. Therefore, a careful balance must be struck to enhance security while minimizing performance degradation. We hope those approaches will inspire future research in developing more efficient and robust FU mechanisms.
@@
We are the first to explore the privacy risks inherent in FU and propose a novel attack method—Federated Unlearning Inversion Attack (FUIA)—which systematically analyzes the security vulnerabilities of FU from the adversarial perspective.
@@
FUIA can effectively target three primary FU approaches: sample unlearning, client unlearning, and class unlearning. The core objective of FUIA is to recover deep information from forgotten data, including feature information or label information, leading to a severe form of privacy leakage.
@@
To mitigate the threats posed by FUIA, we propose two feasible defense strategies: gradient pruning and differential privacy-based noise injection. These defenses provide new insights for designing more secure FU methods in the future.
@@
We conduct comprehensive experiments on multiple benchmark datasets across various FU methods. The results demonstrate that FUIA can effectively recover critical information from forgotten data, highlighting the privacy vulnerabilities of current FU methods.
The rest of this paper is organized as follows. Section II introduces related work. Section III presents the background knowledge of FUIA. Section IV describes the application scenarios and challenges of FUIA. Section V details the implementation of FUIA. Sections VI and VII focus on experimental validation and result analysis. Section VIII discusses potential defense methods. Finally, Section IX concludes the study.
@@ II-A Federated Unlearning
The primary goal of FU is to derive an unlearned model directly from the original model without retraining, thereby significantly improving the efficiency of the unlearning process. Depending on the target data, FU methods can be categorized into sample unlearning, client unlearning, and class unlearning [6]. Existing research primarily focuses on improving the efficiency of FU. For instance, FedU [7] approximates and removes the influence of forgotten data while maintaining model utility for efficient sample unlearning. FRAMU [8] dynamically adjusts data importance using attention mechanisms and federated reinforcement learning, thus realizing sample unlearning. FedEraser [9] leverages historical updates retained by the central server and calibration methods to quickly eliminate client data impact, achieving client unlearning. FedRecovery [10] computes weighted residual gradients of historical updates and applies differential privacy noise to ensure the unlearned model remains statistically indistinguishable from a fully retrained one thereby enabling client unlearning. FURF [12] employs rapid retraining, taylor expansion, and fisher information matrix approximation to efficiently erase designated client data. FUCDP [11] quantifies the contribution of CNN channels to class discrimination using a TF-IDF-based approach and prunes relevant channels to achieve class unlearning. Despite advancements in efficiency, the privacy vulnerabilities of FU remain largely underexplored. Only a few studies have attempted to assess FU-related security risks. For example, FedMUA [14] manipulates the FU process to alter the predictions of the global model, but its primary goal is to disrupt model performance rather than expose potential privacy leaks. Overall, existing research predominantly focuses on optimizing FU efficiency while lacking systematic and in-depth investigations into potential privacy vulnerabilities within the FU process.
@@ II-B Gradient Inversion Attack in Federated Learning
FL is designed to provide default privacy for clients by allowing multiple clients to collaboratively train a model while sharing only their local training parameters with the server, without exposing their raw local data [19]. However, recent studies have shown that the default privacy mechanism in FL is insufficient to prevent training data from being compromised by gradient reconstruction-based privacy leakage attacks. In particular, Gradient Inversion Attacks can infer local training data or related labels by reconstructing gradients, thereby posing a significant threat to client privacy [20].
Existing gradient inversion attacks can be broadly classified into two categories: iterative-based attacks and recursive-based attacks [21]. Iterative-based attacks aim to minimize the distance between virtual gradients and ground-truth gradients. These methods treat the distance between gradients as the error, consider virtual inputs as optimization parameters, and formulate the recovery process as an iterative optimization problem. Representative frameworks include DLG [15], iDLG [22], and STG [23]. On the other hand, recursive-based attacks employ closed-form algorithms to reconstruct the original data. The key insight of these methods lies in leveraging the implicit relationships between input data, model parameters, and per-layer gradients to identify the optimal solution with minimal error. Examples of recursive attacks include R-GAP [24] and COPA [25].
@@ II-C Discussion
Existing research has not specifically investigated the privacy vulnerabilities in the FU process, and there is a lack of dedicated attack methods targeting FU. Inspired by the research on GIA in FL, we propose the FUIA, the first method aimed at studying the privacy vulnerabilities inherent in FU itself. In the current research on machine unlearning, some researchers have proposed the machine unlearning inversion attack (MUIA) [26], but it is not applicable to FL environments. The main reason for this is that in FL training, the interactions in each round cause the mixing of model parameters across clients. This means that the model of each client contains not only parameters from its own data but also parameters from the data of other clients. This mixing of model parameters caused by interactions creates significant challenges for inversion attacks targeting forgotten data on a specific client. Moreover, clients in FL may have highly heterogeneous data characteristics and distributions, which increases the difficulty for inversion attacks to infer specific client data from the global model. Additionally, the aggregation method in FL and other factors can also significantly affect the performance of inversion attacks.
@@ III Preliminaries
And Dist represents a distance function (e.g., distance or cosine similarity), commonly used to measure gradient similarity in GIA. Reg represents a regularization term to ensure realistic sample reconstruction. In image classification tasks, additional regularization techniques such as total variation [28] (to reduce noise) or clipping constraints (to limit pixel values) are often applied to produce natural-looking images. The ultimate goal is to optimize both the virtual samples and virtual labels to retrieve the local training data .
@@ IV Problem Statement
In this section, we present the application scenarios of FUIA, the attack forms, and the associated challenges.
@@ IV-A Threat Model
In this work, we primarily focus on classification tasks, which represent one of the most widely used applications in FL and serve as the foundation for most existing FU studies. We assume that the server acts as an honest-but-curious adversary, meaning that it passively records the parameter updates uploaded by each client throughout the training process and continues to track the parameter updates of target clients (i.e., those submitting unlearn requests) during the unlearning process. Apart from logging these updates, the server does not interfere with the normal training or unlearning procedures. This assumption is both intuitive and subtle, aligning with common FL practices and making detection difficult. In fact, many FU methods inherently rely on storing historical updates to facilitate efficient unlearning [13, 30, 9, 31]. Once the unlearning process is complete, the server leverages the recorded model parameters to analyze differences before and after unlearning, launching an inversion attack to recover deep-level information about the forgotten data. This attack fundamentally compromises the privacy guarantees of FU.
IV-B Problem Formulation and Challenges
@@ V Attack Methodology
In this section, we systematically discuss the specific implementation approaches of FUIA from three perspectives: sample unlearning, client unlearning, and class unlearning.
@@ V-A3 Step 3: Gradient Inversion
Following the previous steps, we extract the gradient information of the target forgotten data from the global model parameters while effectively removing noise from other clients. Inspired by GIA in FL, the goal of gradient inversion is to decode the gradient information of forgotten data using optimization algorithms to recover its features. However, not all GIA optimization algorithms are suitable for FUIA. The primary reason is that, to ensure applicability in a broader range of scenarios and remain inconspicuous, we assume the server, as an attacker, operates under a relatively relaxed setting, behaving as an honest-but-curious adversary rather than a malicious one. Besides, we assume the server does not have explicit knowledge of the FU mechanism. Consequently, the server only obtains an approximate gradient of the forgotten data rather than its exact value. Additionally, some GIA methods require auxiliary information beyond gradients, such as data distribution, to facilitate reconstruction. However, we assume the attacker strictly adheres to the FL protocol and cannot access any additional information beyond the gradients.
To address these challenges, we adopt the optimization algorithm from [33] to recover the feature information of the forgotten data based on its gradient information. The principle of GIA is discussed in Section III. Specifically, we choose cosine similarity as the distance function to measure gradient similarity, focusing more on the direction of the gradients because it contains crucial information on how the model responds to input changes. We also include total variation (TV) as a regularization term to promote smoothness in the reconstructed image while preserving edge information. Based on the steps above, we obtain the gradient information of the target forgotten data on client . Therefore, the objective of the optimization algorithm can be expressed as:
@@
This term helps suppress noise in the reconstructed image while preserving edge details. is a hyperparameter balancing the two terms. The goal is to use the optimization algorithm to make the gradient of the constructed virtual sample as close as possible to the gradient of the target forgotten data, thereby obtaining a virtual sample that closely matches the target forgotten data. Since we have obtained an approximation of the gradient for the forgotten data, we focus more on the direction of the gradients rather than their magnitude, which is why we use cosine similarity as the distance function.

### v3

# Model Inversion Attack against Federated Unlearning

@@ Abstract
With the enactment of regulations related to the “right to be forgotten”, federated learning is facing new privacy compliance requirements. Therefore, federated unlearning (FU) is designed to efficiently remove the influence of specific data from FL models, ensuring that user data can be revoked to enhance privacy. However, existing FU studies primarily focus on improving unlearning efficiency, with little attention given to the potential privacy risks introduced by FU itself. To bridge this research gap, we propose a novel federated unlearning inversion attack (FUIA) to expose potential privacy leakage in FU. This work represents the first systematic study on the privacy vulnerabilities inherent in FU. FUIA can apply to three major FU scenarios: sample unlearning, client unlearning, and class unlearning, demonstrating broad applicability and threat potential. Specifically, the server, acting as an honest-but-curious attacker, continuously records model parameter changes throughout the unlearning process and analyzes the differences before and after unlearning to infer the gradient information of forgotten data, enabling the reconstruction of its features or labels. FUIA directly undermines the goal of FU to eliminate the influence of specific data, exploiting vulnerabilities in the FU process to reconstruct forgotten data, thereby revealing flaws in privacy protection. Moreover, we explore two potential defense strategies that introduce a trade-off between privacy protection and model performance. Extensive experiments on multiple benchmark datasets and various FU methods demonstrate that FUIA effectively reveals private information of forgotten data.
@@
Gradient inversion attack, federated unlearning, machine unlearning, federated learning.I Introduction
With the rapid development of artificial intelligence, the issue of data privacy and security in training machine learning models has garnered increasing attention from both academia and industry [1]. To address these concerns, various regulations have been enacted worldwide, such as the General Data Protection Regulation (GDPR) [2] in the European Union and the California Consumer Privacy Act (CCPA) [3] in the United States. These regulations provide robust legal protections for data owners, with the “right to be forgotten” [4] drawing significant attention in recent years. This right mandates that when a data owner requests the deletion of their data, the model owner must not only remove the corresponding data from the database but also eliminate its impact on the trained model. Federated learning (FL) [5] is designed to enable distributed model training while preserving the privacy of individual clients. However, the introduction of the “right to be forgotten” imposes new requirements on FL, particularly in efficiently fulfilling data owners’ requests for data removal. Recent studies have proposed a novel approach known as federated unlearning (FU) [6], which aims to erase the influence of the target data from the model without the need for retraining the entire model.
@@
During the unlearning process, the server has full access to the parameter differences between the original model and the unlearned model, and it can precisely record the parameter updates uploaded by clients in each training round. However, this extensive access poses significant privacy risks, particularly regarding the leakage of forgotten data. In FL, prior studies have exploited the server’s ability to track model updates and proposed the gradient inversion attack (GIA) [15, 16], which reconstructs clients’ original training data by analyzing parameter changes. Inspired by this, we introduce a novel attack against FU—federated unlearning inversion attack (FUIA), which aims to infer the features or labels information of forgotten data by leveraging parameter changes during the unlearning process. FUIA directly compromises the purpose of FU, as forgotten data remains vulnerable to privacy leakage even after unlearning is completed.
In FUIA, the attacker is modeled as an ”honest-but-curious” server that follows the prescribed training and unlearning protocols while continuously recording client-uploaded parameter updates and analyzing the parameter differences before and after unlearning. This approach of storing parameter updates has been widely adopted in privacy studies of FL and FU, making it a reasonable assumption rather than an additional constraint. As shown in Figure 1, the attacker in FUIA exploits the stored differences between the original and unlearned model parameters to estimate the gradient information of forgotten data, thereby reconstructing its features or inferring its labels information.
@@
Based on different unlearning targets, existing research on FU can be classified into three scenarios: sample unlearning, client unlearning, and class unlearning [17]. Sample unlearning focuses on eliminating the influence of several data samples. Client unlearning requires the model to entirely forget the data from a specific client, making sure that the contributions of the client are no longer reflected in the model parameters. Class unlearning aims to completely remove the influence of a specific class of data from the model. To exploit vulnerabilities in these FU methods, we propose FUIA, which targets all three scenarios of unlearning mechanisms. (1) In the cases of sample unlearning and client unlearning, FUIA aims to recover the features of the forgotten data, such as reconstructing images. The core principle behind this attack lies in the observation that the parameter differences before and after unlearning inherently encode gradient information related to the forgotten data. By leveraging techniques inspired by GIA, which are well-studied in FL, the attacker can apply optimization-based inversion methods to reconstruct the forgotten data. (2) For class unlearning, since the server already possesses knowledge of the global dataset, the goal of FUIA shifts from reconstructing individual data samples to inferring the label of the forgotten class. In general, data from a particular class has a significant impact on model parameters, especially in the output layer. Therefore, the attacker can compare the parameter changes of the output layer before and after unlearning to infer the label of the removed class. The attack strategy of FUIA demonstrates its broad applicability across different FU scenarios.
@@
Moreover, we propose two general defense strategies. First, during the unlearning process in FU, we introduce gradient pruning, where a certain proportion of less important parameter updates from each client is removed before aggregation. The primary goal of this method is to prevent the server from obtaining a complete and precise set of model updates. Second, we leverage differential privacy as an additional defense mechanism. During the unlearning process, we introduce random noise into the model updates submitted by clients, ensuring that the server only receives perturbed parameter updates. These approaches effectively both disrupt the ability of the attackers to infer precise gradient information related to the forgotten data, thereby weakening the effectiveness of FUIA. However, they can lead to poor model performance. Therefore, a careful balance must be struck to enhance security while minimizing performance degradation. We hope those approaches will inspire future research in developing more efficient and robust FU mechanisms.
@@
We are the first to investigate the inherent privacy risks in FU and demonstrate that FUIA can completely undermine the goal of FU by reconstructing the features and labels of forgotten data, leading to severe privacy leakage.
@@
FUIA exhibits a broad attack scope, effectively targeting three primary FU scenarios: sample unlearning, client unlearning, and class unlearning.
@@
We propose and validate two defense strategies that offer new insights to enhance the privacy protection of FU.
@@
We conduct comprehensive experiments on multiple benchmark datasets across three different FU scenarios, demonstrating the effectiveness of FUIA in exposing the privacy information of forgotten data.
@@ II-A Federated Unlearning
The primary goal of FU is to derive an unlearned model directly from the original model without retraining, thereby significantly improving the efficiency of the unlearning process. Depending on the target data, FU methods can be categorized into sample unlearning, client unlearning, and class unlearning [6]. Existing research primarily focuses on improving the efficiency of FU. For instance, FedU [7] approximates and removes the influence of forgotten data while maintaining model utility for efficient sample unlearning. FRAMU [8] dynamically adjusts data importance using attention mechanisms and federated reinforcement learning, thus realizing sample unlearning. FedEraser [9] leverages historical updates retained by the central server and calibration methods to quickly eliminate client data impact, achieving client unlearning. FedRecovery [10] computes weighted residual gradients of historical updates and applies differential privacy noise to ensure the unlearned model remains statistically indistinguishable from a fully retrained one, thereby enabling client unlearning. FURF [12] employs rapid retraining, taylor expansion, and fisher information matrix approximation to efficiently erase designated client data. FUCDP [11] quantifies the contribution of CNN channels to class discrimination using a TF-IDF-based approach and prunes relevant channels to achieve class unlearning.
@@ II-B Gradient Inversion Attack in Federated Learning
FL is designed to provide default privacy for clients by allowing multiple clients to collaboratively train a model while sharing only their local training parameters with the server without exposing their raw local data [18]. However, recent studies have shown that the default privacy mechanism in FL is insufficient to prevent training data from being compromised by gradient reconstruction-based privacy leakage attacks. In particular, GIA can infer local training data by reconstructing gradients, thereby posing a significant threat to client privacy [19]. Existing gradient inversion attacks can be broadly classified into two classes: iterative-based attacks and recursive-based attacks [20]. Iterative-based attacks aim to minimize the distance between virtual gradients and ground-truth gradients. These methods treat the distance between gradients as the error, consider virtual inputs as optimization parameters, and formulate the recovery process as an iterative optimization problem. Representative frameworks include DLG [15], iDLG [21], and STG [22]. On the other hand, recursive-based attacks employ closed-form algorithms to reconstruct the original data. The key insight of these methods lies in leveraging the implicit relationships between input data, model parameters, and per-layer gradients to identify the optimal solution with minimal error. Examples of recursive attacks include R-GAP [23] and COPA [24].
@@ II-C Discussion
Existing research predominantly focuses on optimizing FU efficiency while lacking systematic and in-depth investigations into potential privacy vulnerabilities within the FU process. Only a few studies have attempted to assess its security risks. For example, FedMUA [14] manipulates the FU process to alter the predictions of the global model, but its primary goal is to disrupt model performance rather than expose potential privacy leaks. Moreover, in the current research on machine unlearning, some researchers have proposed the machine unlearning inversion attack (MUIA) [25], but it is not applicable to FL environments. The main reason for this is that in FL training, the interactions in each round cause the mixing of model parameters across clients. This means that the model of each client contains not only parameters from its own data but also parameters from the data of other clients. This mixing of model parameters caused by interactions creates significant challenges for inversion attacks targeting forgotten data on a specific client. Moreover, clients in FL may have highly heterogeneous data characteristics and distributions, which increases the difficulty for inversion attacks to infer specific client data from the global model. And the aggregation method in FL and other factors can also significantly affect the performance of inversion attacks. Therefore, to bridge this research gap, inspired by the study of GIA in FL, we propose FUIA. This is the first targeted investigation into potential privacy vulnerabilities inherent in FU. Compared to existing research, FUIA exhibits significant differences and holds great significance for exploring more secure FU mechanisms.
@@ III Preliminaries
And Dist(·) represents a distance function (e.g., distance or cosine similarity), commonly used to measure gradient similarity in GIA. Reg(·) represents a regularization term to ensure realistic sample reconstruction. In image classification tasks, additional regularization techniques such as total variation [27] (to reduce noise) or clipping constraints (to limit pixel values) are often applied to produce natural-looking images. The ultimate goal is to optimize both the virtual samples and virtual labels to retrieve the local training data .
@@ IV-A Threat Model
In this study, we primarily focus on classification tasks, which represent one of the most widely used applications in FL and serve as the foundation for most existing FU studies. We assume that the server acts as an honest-but-curious attacker, meaning that it passively records the parameter updates uploaded by each client throughout the training process and continues to track the parameter updates of target clients (i.e., those submitting unlearning requests) during the unlearning process. Apart from logging these updates, the server does not interfere with the normal training or unlearning procedures. This assumption is both intuitive and subtle, aligning with common FL practices and making detection difficult. In fact, many FU methods inherently rely on storing historical updates to facilitate efficient unlearning [13, 29, 9, 30]. Once the unlearning process is complete, the server leverages the recorded model parameters to analyze differences before and after unlearning, launching an inversion attack to recover deep-level information about the forgotten data. This attack fundamentally compromises the privacy guarantees of FU.
IV-B Problem Formulation
@@ V-A3 Step 3: Gradient Inversion
Following the previous steps, we extract the gradient information of the target forgotten data from the global model parameters while effectively removing noise from other clients. Inspired by GIA in FL, the goal of gradient inversion is to decode the gradient information of forgotten data using optimization algorithms to recover its features. However, not all GIA optimization algorithms are suitable for FUIA. The primary reason is that, to ensure applicability in a broader range of scenarios and remain inconspicuous, we assume the server, as an attacker, operates under a relatively relaxed setting, behaving as an honest-but-curious attacker rather than a malicious one. Besides, we assume the server does not have explicit knowledge of the FU mechanism. Consequently, the server only obtains an approximate gradient of the forgotten data rather than its exact value. Additionally, some GIA methods require auxiliary information beyond gradients, such as data distribution, to facilitate reconstruction. However, we assume the attacker strictly adheres to the FL protocol and cannot access any additional information beyond the gradients.
To address these challenges, we adopt the optimization algorithm from [32] to recover the feature information of the forgotten data based on its gradient information. The principle of GIA is discussed in Section III. Specifically, we choose cosine similarity as the distance function to measure gradient similarity, focusing more on the direction of the gradients because it contains crucial information on how the model responds to input changes. Besides, since we have obtained an approximation of the gradient for the forgotten data instead of exact values, we focus more on the direction of the gradients rather than their magnitude. We also include total variation (TV) as a regularization term to promote smoothness in the reconstructed image while preserving edge information. Based on the steps above, we obtain the gradient information of the target forgotten data on client . Therefore, the objective of the optimization algorithm can be expressed as:
@@
This term helps suppress noise in the reconstructed image while preserving edge details. is a hyperparameter balancing the two terms. The goal is to use the optimization algorithm to make the gradient of the constructed virtual sample as close as possible to the gradient of the target forgotten data, thereby obtaining a virtual sample that closely matches the target forgotten data. Additionally, during the optimization process, we assume that the data labels are known. As mentioned earlier, previous studies have shown that data labels can be directly extracted in GIA [22, 28].